A body moves with a velocity of 3m/s due east and then turns due  north to travel with the same velocity.If the total time of travel is 6s,the acceleration of the body is

Grade:7

1 Answers

Arun
5 years ago
Acceleration = resultant velocity ÷ total time 
Since these are perpendicular velocities the resultant velocity will be the hypotenuse. 
Hypotenuse = √(3²+3²) = √18 = 4.24m/s
Acceleration = 4.24/6 = 0.71m/s²